OpenStudy (lemanual):

Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. Equation 1 Equation 2 |5x − 6| = −41 |7x + 13| = 27 Equation 1 has more solutions than equation 2. Equation 1 and Equation 2 have the same number of solutions. Equation 2 has more solutions than Equation 1. The number of solutions cannot be determined.

OpenStudy (anonymous):

c

OpenStudy (lemanual):

Thanks

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do want to know why

OpenStudy (lemanual):

Yes :D

OpenStudy (anonymous):

ok. Equation 1 have no solution

OpenStudy (texaschic101):

| 5x - 6 | = - 41 5x - 6 = -41 -(5x - 6) = - 41 5x = -41 + 6 -5x + 6 = -41 5x = - 35 -5x = -41 - 6 x = - 7 -5x = -47 x = 47/5 two solutions | 7x + 13 | = 27 7x + 13 = 27 - (7x + 13) = 27 7x = 27 - 13 -7x - 13 = 27 7x = 14 -7x = 27 + 13 x = 2 -7x = 40 x = -40/7 two solutions answer is B

OpenStudy (anonymous):

and Equation 2 is 2,-40/7

OpenStudy (lemanual):

Ok Thanks :D